OpenStudy (maddytarin_98):

Micaela is 2 years older than Sam. In 4 years, the sum of their ages will be 40. How old is Micaela now? Describe and fix any errors in the solution. Let m = Micaela’s age and s = Sam’s age. m = s + 2 (s + 4) + (s + 2 + 4) = 40 s = 15 Sam is 15.

OpenStudy (maddytarin_98):

someone please help me

OpenStudy (maddytarin_98):

whats the mistake in the problem?

OpenStudy (maddytarin_98):

because i have to write a response and i have to mention key words that the system will read

OpenStudy (photon336):

m = s+6 sam s+4 2s+10 = 40 s = 15 sam is 15 and Michela is 17 4 years from now 15+4 = 19 for sam 21 for michela 19+21 = 40 years.

OpenStudy (photon336):

so yeah I looked at this again and you're right sam is 15 years old now

OpenStudy (photon336):

but remember Micaela is 2 years older than sam?

OpenStudy (maddytarin_98):

yeah!

OpenStudy (maddytarin_98):

thank you!

geerky42 (geerky42):

Nothings wrong with solution?